OpenStudy (anonymous):

Solve the following. One car left Carville driving toward Beattyville which is 300 miles away. A second car left Beattyville at the same time, driving toward Carville. If the first and second cars are averaging 55 mph and 65 mph, in how many hours will they pass each other?

OpenStudy (anonymous):

RT + RT = D 55x + 65x = 300 120x = 300 x = 2.5 They will pass each other in 2.5 hours. <==ANSWER Two jets are beginning their flight at the Louisville International Airport. One starts at 1:00 p.m. and flies east at 395 mph. Another starts one hour later and flies west at 400 mph. In how many hours after the first plane left, will they be 2,000 miles apart? Round to the nearest tenth of an hour. RT + R(T - 1) = D 395x + 400(x - 1) = 2000 395x + 400x - 400 = 2000 795x = 2400 x = 3.018 hours Rounded off, it will be 3.0 hours after the first plane takes off. <==ANSWER Harold left town driving east at 55 mph. Carol left 2 hours later driving the same route at 65 mph. How many hours after Harold left will it take Carol to catch up to Harold? RT = R(T - 2) 55x = 65(x - 2) 55x = 65x - 130 -10x = -130 x = 13 It will be 13 hours after Harold left until Carol catches up with him. <==ANSWER